Three siblings have a combines age of 108 yrs. The oldest is 8 years older than the youngest. what are the ages of the siblings?
Answer by ankor@dixie-net.com(22740)

Three unknowns, two equations, calls for some reasoning here.
Three siblings have a combined age of 108 yrs.
a + b + c = 108
;
The oldest is 8 years older than the youngest.
a = c + 8
:
a + b + c = 108
(c+8) + b + c = 108
2c + b = 108 - 8
2c = b = 100
b = -2c + 100
we know b is < 8 and > 1 yr more than c
c = 31
b = -2(31) + 100
b = 38
or
c = 32
b = -2(32) + 100
b = 36
and
c = 33
b = -2(33) + 100
b = 34
:
So we have three possibilities:
a = 39
b = 38
c = 31
or
a = 40
b = 36
c = 32
or
a = 41
b = 34
c = 33